# DoorINet: A Deep-Learning Inertial Framework for Door-Mounted IoT Applications

### v1

@@ Abstract
Many Internet of Things applications utilize low-cost, micro, electro-mechanical inertial sensors. A common task is orientation estimation. To tackle such a task, attitude and heading reference system algorithms are applied. Relying on the gyroscope readings, the accelerometer readings are used to update the attitude angles, and magnetometer measurements are utilized to update the heading angle. In indoor environments, magnetometers suffer from interference that degrades their performance. This mainly influences applications focused on estimating the heading angle like finding the heading angle of a closet or fridge door. To circumvent such situations, we propose DoorINet, an end-to-end deep-learning framework to calculate the heading angle from door-mounted, low-cost inertial sensors without using magnetometers. To evaluate our approach, we record a unique dataset containing 391 minutes of accelerometer and gyroscope measurements and corresponding ground-truth heading angle. We show that our proposed approach outperforms commonly used, model based approaches and data-driven methods. To enable reproducibility of our results and future research, both code and data are available at https://github.com/ansfl/DoorINet.
@@ I Introduction
The Internet of Things (IoT) is essentially a network of devices that can collect information and connect to and communicate with each other without human interaction. Remote data storage and cloud computing allow automatic data analysis and decision-making capabilities. IoT-enabled devices combine multiple sensors such as inertial sensors, and can be a part of various applications such as motion capture and human activity recognition [1, 2, 3, 4], , sports applications [5, 6, 7, 8], health-related applications [9, 10, 11], pedestrian dead reckoning [12, 13, 14], and smart home applications [15, 16, 17].
@@
Inertial measurement units (IMU) consist of three mutually orthogonal accelerometers and three gyroscopes aligned with the accelerometers. Additionally, there can be three orthogonal magnetometers aligned in the same direction as other sensors. Accelerometers provide the specific force vector, gyroscopes the angular rate vector, and the magnetometers the magnetic field vector. IMUs come in a wide range of performance grades intended for different applications. Though there is no universally agreed definition of high-, medium-, and low-grade IMUs, they can be grouped into several categories: marine, aviation, intermediate, tactical, and automotive [18]. Automotive grade is sometimes called consumer or commercial grade. The most affordable IMU sensors are usually built with micro electro-mechanical sensor (MEMS) technology, also making them small compared to mechanical IMUs. Miniaturized MEMS IMUs are widely used in consumer products, i.e., smartphones, IoT applications, and wearable devices.
@@
In non-navigational applications (such as most IoT applications) inertial sensor readings are used as input to attitude and heading reference system (AHRS) algorithms that provide the orientation of a device; that is, the three Euler angles with respect to a fixed coordinate frame, which uniquely determine the orientation.
@@
In model-based approaches the AHRS algorithm divides into two separate independent parts: orientation propagation from a gyroscope and readings followed by updates from accelerometer and magnetometer measurements. In nonlinear filtering approaches the attitude kinematic equations are treated as the system model, and the accelerometer and magnetometer observations are processed in the measurement model to update the state and propagate the error-state covariance matrix [19]. Another commonly used approach is complementary filtering, which combines compensated gyroscope measurements with gain-multiplied measurements from accelerometers and magnetometers [20, 21]
@@
In learning-based approaches, IMU readings are fed directly into the deep-learning algorithm to provide attitude and heading estimation [22]. Alternatively, a deep-learning algorithm can be used for denoising inertial measurements, and attitude estimation is obtained using an integrated processed ”noise-free” signal like in [23]. Hybrid approaches try to combine the best of model and learning approaches. For example, the deep attitude estimator [24, 25, 26] uses a deep-learning model to calculate gain for the complementary filter algorithm structure.
Indoor environments are full of electromagnetic fields, resulting in constant and time varying influence on magnetometer readings. This is even more relevant in specific industrial buildings with high levels of magnetic field where the heading angle estimation accuracy degrades.
@@
In this paper we introduce and address the problem of door mounted IMUs. Here, the goal is to determine the opening angle of a door. This information may be utilized in several IoT applications such as smart homes, smart offices, or building management. To cope with low magnetometer performance in an indoor enthronement, we propose DoorINet, a deep-learning, end-to-end framework for estimating the heading angle of door-mounted IMU using only accelerometers and gyroscopes. Two versions of DoorINet are examined: AG-DoorINet that takes accelerometer and gyroscope measurements, and G-DoorINet that takes only gyroscope readings.
@@
To evaluate our proposed approaches we re-record a unique dataset using Xsens DOT IMUs mounted on three different doors, resulting in 391 minutes of recorded inertial data and accurate headings obtained from a higher grade Memsense IMU. We compare our approach with two model-based and three learning-based AHRS approaches and show that ours outperform the rest on a real-life scenario of an inner door over 90 minutes. Our deep-learning framework is able to generalize over different error parameters of different IMU sensors, and it is able to correctly calculate the heading angle.
@@
The rest of the paper is organized as follows: Section II introduces model-based approaches for heading estimation. Section III describes our proposed learning approach for heading determination. Section IV explains our dataset collection setup while Section V describes post-processing procedures. Section VI shows the experimental results and Section VII concludes this work.
II Model-Based AHRS Algorithms
@@ II-A Direct Integration
where $\psi_{0}$ is the initial heading angle. This approach does not take into account error factors in the gyroscope sensor and alignment errors and therefore is considered less accurate than AHRS methods.
II-B Madgwick Filter
Madgwick first proposed an orientation filter for inertial and inertial/magnetic sensor arrays in [27], and later its extended version [28]. We follow a version of the Madgwick algorithm as presented in chapter 7 of [20]. This algorithm functions as a complementary filter that combines compensated gyroscope measurements with filtered measurements from the accelerometer with a frequency determined by the gain.
@@ II-B Madgwick Filter
III Proposed Approach
@@
Figure 1: Proposed data-driven approach.
As IoT applications are considered, we focused our design on relatively shallow architectures allowing their implementation for real-time settings. Two types of networks are proposed, differing in their structure and input, as addressed in the next section.
@@ III-A Gyroscope-Only DoorINet Architecture (G-DoorINet)
Bidirectional recurrent layers were introduced in [29] and they connect hidden states of two recurrent layers processing an input sequence in both forward and backward directions to the same output.
@@
In addition to BiGRU layers we employ fully connected (FC) layers, which were first introduced in [31], and calculate vector-matrix multiplication $\boldsymbol{\mathrm{x}}\times\boldsymbol{\mathrm{W}}$, where $\boldsymbol{\mathrm{x}}$ is the input vector and $\boldsymbol{\mathrm{W}}$ is the weight matrix. An activation function tanh is applied between fully connected layers element-wise and defined as
@@
Our gyroscope-only DoorINet (G-DoorINet) architecture consists of consequent bi-directional GRU (BiGRU) layers that feed on gyroscope measurements and then FC layers as presented in Figure 2:
@@ IV-A2 Session 2
In this session three experiments were recorded according to the Setup 2. An additional experiment was recorded in a distant location from the door hinge with IMUs rotated 90°clockwise. At each location two experiments were recorded with a total of eight experiments. In each experiment the door was opened to a predefined angle and closed shut five times. The approximate chosen angles for the experiments were 32°and 64°.
@@ IV-A3 Session 3
In this session eight experiments were recorded in total. In each, the door was opened to a predefined angle and closed shut, 10-11 times. In the last three experiments the door was opened to 90° at different speeds: relatively slow, medium, and fast.
@@ IV-C Dataset summary
V Data processing, model implementation, and training
V-A Preprocessing steps: train dataset
@@
V-B Preprocessing steps: test dataset
For the GT heading angle (generated by Memsense IMU3025 IMU) a modified ”thresholded” version of Madgwick algorithm was introduced. At each time frame it compares the gyroscope vector norm with a predefined threshold, and if the vector norm is below the threshold, the algorithm assumes that there is no movement at that specific time frame and keeps the last value of the heading angle. As a consequence, the GT heading was altered to reflect the actual conditions of the shut door periods as shown in Figure 9.
@@ V-B Preprocessing steps: test dataset
V-C Model implementation and training
@@
VI Experimental results
VI-A Metrics
@@
VI-B Competitive approaches
There exist a significant number of data-driven models for orientation estimation. Most of them solve a 3D estimation problem, which in our case is superficial as we are focused only on the heading angle estimation. We compare our approaches to five other methods. Two are model-based approaches: the direct integration approach, as described in Section II-A, and the Madgwick filter, presented in Section II-B. In addition, three data-driven approaches were implemented for comparison with our approach:
VI-B1 Deep Attitude Estimator (DAE) [25]
VI-B2 Quaternion Model A [22]
VI-B3 LGC-Net [23]
VI-C Results
As described in Table III, the test dataset includes recordings of DOT #12 that were not used in the train and validation dataset, IMU #14 that was used in Session 1 to generate 7.2% of the data used for training and validation, and IMU #5 that was used in Session 3 to generate 11.5% of the training/validation data sets. The following figures presenting the results are divided into two parts. The top plot shows all of the examined model outputs and heading GT, while the bottom plot shows the heading error of each model.
@@ VI-C Results
Figure 12 presents the estimated heading angle of our two approaches and five other approaches as described in the previous section. The heading was estimated using the test dataset of DOT #12, which was not present in the train dataset. As presented in the Figure 12, all of the tested algorithms perform worse than the proposed DoorINet models, and their performance becomes worse with time (the difference becomes bigger). The figure also shows that the Quaternion ModelA fails to calculate the correct angle for the door opening. Both of our approaches obtained the best performance where AG-DoorINet has a minimum RMSE of 1.24 degrees.
@@ VII Conclusion
DoorINet is suitable for different smart home and office applications; for example, smart office, home office, or building management.
In addition, DoorINet can be implemented on any other AHRS application. In future work, we aim to extend DoorINet to include estimation of the roll and pitch angles.

### v2

# DoorINet: Door Heading Prediction through Inertial Deep Learning

@@ Abstract
Inertial sensors are widely used in a variety of applications. A common task is orientation estimation. To tackle such a task, attitude and heading reference system algorithms are applied. Relying on the gyroscope readings, the accelerometer measurements are used to update the attitude angles, and magnetometer measurements are utilized to update the heading angle. In indoor environments, magnetometers suffer from interference that degrades their performance resulting in poor heading angle estimation. Therefore, applications that estimate the heading angle of moving objects, such as walking pedestrians, closets, and refrigerators, are prone to error. To circumvent such situations, we propose DoorINet, an end-to-end deep-learning framework to calculate the heading angle from door-mounted, low-cost inertial sensors without using magnetometers. To evaluate our approach, we record a unique dataset containing 391 minutes of accelerometer and gyroscope measurements and corresponding ground-truth heading angle. We show that our proposed approach outperforms commonly used, model based approaches and data-driven methods. To enable reproducibility of our results and future research, both code and data are available at https://github.com/ansfl/DoorINet.
@@ I Introduction
Inertial sensors are found in many applications such as motion capture and human activity recognition [1, 2, 3, 4], sports applications [5, 6, 7, 8], health-related applications [9, 10, 11], pedestrian dead reckoning [12, 13, 14], and smart home applications [15, 16, 17].
@@
Inertial measurement units (IMU) consist of three mutually orthogonal accelerometers and three gyroscopes, aligned with the accelerometers. Additionally, there can be three orthogonal magnetometers, aligned in the same direction as other sensors. Accelerometers provide the specific force vector, gyroscopes - the angular rate vector, and the magnetometers - the magnetic field vector. IMUs come in a wide range of performance grades intended for different applications. Though there is no universally agreed definition of high-, medium-, and low-grade IMUs, they can be grouped into several categories: marine, aviation, intermediate, tactical, and automotive [18]. Automotive grade is sometimes called consumer or commercial grade. The most affordable IMU sensors are usually built with micro electro-mechanical sensor (MEMS) technology, also making them small compared to mechanical IMUs. Miniaturized MEMS IMUs are widely used in consumer products, i.e., smartphones, IoT applications, and wearable devices.
@@
In non-navigational applications (such as most IoT applications), inertial sensor readings are used as input to attitude and heading reference system (AHRS) algorithms that provide the orientation of a device; that is, the three Euler angles with respect to a fixed coordinate frame, which uniquely determine the orientation.
@@
In model-based approaches, the AHRS algorithm divides into two separate independent parts: orientation propagation from a gyroscope, and readings followed by updates from accelerometer and magnetometer measurements. In nonlinear filtering approaches the attitude kinematic equations are treated as the system model, and the accelerometer and magnetometer observations are processed in the measurement model to update the state and propagate the error-state covariance matrix [19]. Another commonly used approach is complementary filtering, which combines compensated gyroscope measurements with gain-multiplied measurements from accelerometers and magnetometers [20, 21]
@@
In learning-based approaches, IMU readings are fed directly into the deep-learning algorithm to provide attitude and heading estimation [22]. Alternatively, a deep-learning algorithm can be used for denoising inertial measurements, and attitude estimation is obtained using an integrated processed ”noise-free” signal like in [23].
Hybrid approaches try to combine the best of model and learning approaches. For example, the deep attitude estimator [24, 25, 26] uses a deep-learning model to calculate gain for the complementary filter algorithm structure.
@@
The heading information is critical for several IoT applications such as smart homes, smart offices, or building management. It is also important for space planning, furniture placement and safety. The heading of a door can be estimated using magnetometers. Yet, indoor environments are full of electromagnetic fields, resulting in constant and time varying influence on magnetometer readings. This is even more relevant in specific industrial buildings with high levels of magnetic field, where the heading angle accuracy degrades. Consequently, magnetometers cannot be used to determine the door heading angle. To cope with this situation we propose DoorINet, a deep learning end-to-end framework for estimating the heading angle of door-mounted IMU using only accelerometers and gyroscopes. Two versions of DoorINet are examined: AG-DoorINet that takes accelerometer and gyroscope measurements, and G-DoorINet that takes only gyroscope readings.
@@
To evaluate our proposed approaches we rrecord a unique dataset using Xsens DOT IMUs mounted on three different doors, resulting in 391 minutes of recorded inertial data and accurate headings obtained from a higher grade Memsense IMU. We compare our approach with two model-based and three learning-based AHRS approaches, and show that ours outperform the rest on a real-life scenario of an inner door over 90 minutes. Our deep-learning framework is able to generalize over different error parameters of different IMU sensors, and it is able to correctly calculate the heading angle.
@@
The rest of the paper is organized as follows: Section II introduces model-based approaches for heading estimation. Section III describes our proposed learning approach for heading determination. Section IV explains our dataset collection setup and describes post-processing procedures. Section V shows the experimental results and Section VI concludes this work.
II Heading Estimation Approaches
II-A Model-based approaches
@@ II-A1 Direct Integration
where $\psi_{0}$ is the initial heading angle. This approach does not take into account error factors in the gyroscope sensor and alignment errors, and, therefore, is considered less accurate than AHRS methods.
II-A2 Madgwick Filter
Madgwick first proposed an orientation filter for inertial and inertial/magnetic sensor arrays in [27], and later its extended version [28]. We follow a version of the Madgwick algorithm as presented in chapter 7 of [20]. This algorithm functions as a complementary filter that combines compensated gyroscope measurements with filtered measurements from the accelerometer with frequency determined by the gain.
@@ II-A2 Madgwick Filter
II-B Learning approaches
There exist a significant number of data-driven models for orientation estimation. Most of them solve a 3D estimation problem, which in our case is superficial as we are focused only on the heading angle estimation. Three data-driven approaches were implemented for comparison with our proposed approach:
II-B1 Deep Attitude Estimator (DAE) [25]
II-B2 Quaternion Model A [22]
II-B3 LGC-Net [23]
III Proposed Learning Approach for Heading Estimation
@@
Figure 1: Our proposed data-driven approach.
As IoT applications are considered, we focused our design on relatively shallow architectures allowing their implementation for real-time settings. Two types of networks are proposed, differing in their structure and input, as addressed in the next section. To find the most suitable network for heading estimation, we examined many different network architectures, including convolutional neural networks (CNN), long-short-term memory layers (LSTM), and gated recurrent networks (GRU). Of all these architectures, GRU was found the most promising and is described in the following sections.
@@ III-A Gyroscope-Only DoorINet Architecture (G-DoorINet)
Bidirectional recurrent layers were introduced in [31], and they connect hidden states of two recurrent layers processing an input sequence in both forward and backward directions to the same output.
@@
In addition to BiGRU layers, we employ fully connected (FC) layers, which were first introduced in [33], and calculate vector-matrix multiplication $\boldsymbol{\mathrm{x}}\times\boldsymbol{\mathrm{W}}$, where $\boldsymbol{\mathrm{x}}$ is the input vector and $\boldsymbol{\mathrm{W}}$ is the weight matrix. An activation function tanh is applied between fully connected layers element-wise and defined as
@@
Figure 2 describes DoorINet (G-DoorINet) architecture for the learning the door heading angle. It takes gyroscope readings as its input and fed them into the bi-directional GRU layer (marked as BiGRU1 box), which output then goes to another bi-directional GRU layer (BiGRU2). Next, the output of BiGRU2 goes to a sequence of eight fully connected layers, which outputs the heading angle:
@@
The BiGRU layer that processes gyroscope readings (BiGRU1) accumulates knowledge about every sample (which is the sequence of 20 consequent readings) in one block of data. Because each ground-truth value is the angle difference between the first and the last heading angle of the IMU, the direction of data is not relevant, therefore use of bidirectional option of GRU layers is justified. The BiGRU2 deals with processed data from gyroscope. Fully connected layers are organized in common pyramid-like encoder fashion and serve as a function approximator that outputs one value for the series of readings.
@@ III-B Accelerometer and Gyroscope DoorINet Architecture (AG-DoorINet)
The BiGRU layers that independently process accelerometer and gyroscope readings (BiGRU1 and BiGRU2) accumulate knowledge about every sample (which is the sequence of 20 consequent readings) in one block of data. Because each ground-truth value is the angle difference between the first and the last heading angle of the IMU, the direction of data is not relevant, therefore use of bidirectional option of GRU layers is justified. The BiGRU3 deals with the stacked processed data from accelerometer and gyroscope, taking it to the next logical layer. Fully connected layers are organized in common pyramid-like encoder fashion and serve as a function approximator that outputs one value for the series of readings.
@@ IV-A2 Session 2
In this session, three experiments were recorded according to the Setup 2. An additional experiment was recorded in a distant location from the door hinge with IMUs rotated 90°clockwise. At each location two experiments were recorded with a total of eight experiments. In each experiment the door was opened to a predefined angle and closed shut five times. The approximate chosen angles for the experiments were 32°and 64°.
@@ IV-A3 Session 3
In this session eight experiments were recorded in total. In each, the door was opened to a predefined angle and closed shut 10-11 times. In the last three experiments the door was opened to 90° at different speeds: relatively slow, medium, and fast.
@@ IV-B Test dataset
The GT heading angle was generated using the Memsense IMU3025 IMU readings. It was placed next to the low-cost IMUs that are the units under test.
@@ IV-C Dataset summary
IV-D Preprocessing steps: train dataset
@@
IV-E Preprocessing steps: test dataset
A modified ”thresholded” version of Madgwick algorithm was introduced to calculate GT heading angle from these IMU readings. At each time frame it compares the gyroscope vector norm with a predefined threshold, and, if the vector norm is below the threshold, the algorithm assumes that there is no movement at that specific time frame, and keeps the last value of the heading angle. As a consequence, the GT heading was altered to reflect the actual conditions of the shut door periods as shown in Figure 9.
@@ IV-E Preprocessing steps: test dataset
IV-F Model implementation and training
@@
V Experimental results
V-A Metrics
@@
V-B Results
As described in Table III, the test dataset includes recordings of DOT #12 that were not used in the train and validation dataset, IMU #14 that was used in Session 1 to generate 7.2% of the data used for training and validation, and IMU #5 that was used in Session 3 to generate 11.5% of the training/validation data sets.
@@ V-B Results
Figures 12-14 present the heading estimation versus time for three different IMUs applied on the test dataset. The heading estimation approaches include our two DoorINet architectures and five other approaches as described in Section II.
As presented in the Figure 12, all of the tested algorithms perform worse than the proposed DoorINet models, and their performance becomes worse with time (the difference becomes bigger). The figure also shows that the Quaternion ModelA fails to calculate the correct angle for the door opening. Both of our approaches obtained the best performance where AG-DoorINet has a minimum RMSE of 1.24 degrees.
@@ VI Conclusion
DoorINet is suitable for different smart home and office applications; for example, smart office, home office, or building management. In addition, DoorINet can be implemented on any other AHRS application. In future work, we aim to extend DoorINet to include estimation of the roll and pitch angles.